Determine the coefficent of the term containing a^5 in the expansion of
(a - 1/a)^5

I got 1 for the coefficent of a^5. However, I can not read my teachers writing and I think that the the question could be, "Determine the coefficent of the term containing a in the expansion of (a - 1/a)^5". If this is the case I am not sure of the coefficent since there are two a's and when you expand using binomial expansion, there are two possible terms that have a values (either multipy the a's together to get a or there is just an a in one of the terms).
Thank you,
Kristen

Answer by stanbon(75887)

Determine the coefficent of the term containing a^5 in the expansion of
(a - 1/a)^5
= (a - a^-1)^5
1st term of the expansion is 5C5a^5(a^-1)^0 = a^5
==========
Cheers,
Stan H.